OpenStudy (anonymous):

Help? 1. If f(x) = |x + 6|, find f(−10). A. -16 B. 16 C. -4 D. 4 2. If f(x) = |x − 4| + 5, find f(1). A. 8 B. 2 C. 0 D. 10 3. If f(x) = 8 − 3|x + 2|, find f(−6). A. 20 B. -16 C. -4 D. 4

OpenStudy (anonymous):

@hwtrain Help plz?

OpenStudy (anonymous):

The absolute value operator essentially makes whatever is inside the vertical lines positive, regardless of its sign. For 1, we have f(-10) of f(x) = |x+6|, so we plug x=-10 into |x+6| to yield |-10+6| = |-4|. Because the absolute value turns everything inside it positive, the answer to question 1 is D) 4. We can take similar steps for question 2. |x-4| for x=1 is |1-4| = |-3| = 3, then add 5 to that to get A) 8. For question 3, |x+2| for x=-6 is |-6+2|=|-4|=4, then multiply by 3 to get 12, then subtract that from 8 to get C) -4.